OpenStudy (anonymous):

If the dimensions of a rectangle become one-third of the original dimensions, then area of the new rectangle will be _______of the area of the original rectangle.

OpenStudy (campbell_st):

its the linear ratio squared so (1/3)^2 = 1/9 the original

OpenStudy (campbell_st):

you can test this by a reactangle 3 x 6 = 18 1/3 measurements 1 x 2 = 2 2/18 = 1/9